# Extreme events in a random set of nonlinear elastic bending waves

### v1

@@ III.2 Correlation
This study leverages the advantages of bending elastic waves, where we observe extreme states similar to those in surface gravity waves but driven by different mechanisms. By using forcing frequencies nearly 100 times higher than those in surface gravity wave experiments, we can collect much more statistical data in a shorter time. These features provide new insights into the mechanisms behind bending elastic waves, offering a different perspective. Finally, we acknowledge that our current definition of local wave period may not fully capture true rogue waves in a set of bending waves, a point to be addressed once the definition is clarified.
@@ Acknowledgements.
We wish to acknowledge the ANR for financial support and B. Gallet, C. Touzé, S. Varghese for fruitful discussions.
Appendixes
@@ Appendix A Tables
$\eta_{\text{rms}(mm)}$00footnotetext: $\eta_{\text{rms}}$ is the r.m.s transversal displacement; $1.159$ $1.158$ $2.230$ $3.596$
$V_{\text{rms,inj}(\frac{m}{s})}$ 00footnotetext: $V_{\text{rms,inj}}$ is the r.m.s injected velocity; $0.179$ $0.180$ $0.312$ $0.464$
K00footnotetext: K is the kurtosis; S is the skewness; $3.057$ $3.010$ $3.105$ $3.057$
@@
$N_{tot}$00footnotetext: $N_{tot}$ is the total number of all recorded events; $5.484\times 10^{5}$ $3.454\times 10^{6}$ $5.046\times 10^{5}$ $4.139\times 10^{5}$
$N_{re}$ 00footnotetext: $N_{re}$ is the number of detected rare events; $162$ $989$ $308$ $160$
frac 00footnotetext: frac is the ratio of $\frac{N_{re}}{N_{tot}}$; $0.00029$ $0.00028$ $0.00061$ $0.00038$

### v2

@@ IV Discussion &\&& Conclusions
A fundamental difference comes from the spectra of both wave fields. Usually, the sea spectrum is much narrower than the spectra plotted in Figure 4. The fundamental modes of the elastic plate play a major role, and the rare events correspond to the longest wavelength according to our statistical definition of the extreme events. This correlation has not been reported for surface waves. In contrast, the rogue waves occur near the Tayfun period $1/f_{T}$, where the Tayfun frequency is $f_{T}=\tilde{f_{o}}/\left[1+(\Delta f/f_{o})^{2}\left(1+(\Delta f/f_{o})^{2}\right)^{-3/2}\right]$[23, 2, 6]. This raises a question: in our setup, there is no inverse cascade, yet energy equipartition between modes at low wavelength is enough to populate the fundamental mode of the plate. In surface waves, where an inverse cascade should enhance the lowest mode, the inverse cascade shrinks to a small range. What are the dissipative phenomena that shrink the inverse cascade and limit the growth of the first mode in wave basins?
This study leverages the advantages of bending elastic waves, where we observe extreme states with an occurrence similar to those in surface gravity waves but driven by different mechanisms. By using forcing frequencies nearly 100 times higher than those in surface gravity wave experiments, we can collect much more statistical data in a shorter time. This allows us to get statistical properties of the wave conditioned by the presence of extreme events. This provides valuable input for understanding the underlying mechanisms driving these extreme events.
One main drawback is that these measurements are limited to single-point data. Although our study provides valuable insight, wave phenomena are inherently spatio-temporal. Point measurements lack the spatial resolution needed to understand many features like energy redistribution or wavefront evolution. Hence, our interpretations of the correlations might fail to capture the complete picture. There are many spatio-temporal measurement techniques that could be applied on the plate to capture full wave dynamics, such as the FTP method [21]. However, simply recording indefinitely would require an immense amount of data storage and processing power, let alone successfully capturing the moment when rare events occur. Using event-based cameras or scanning laser vibrometers could offer possible solutions.
@@ Acknowledgements.
We wish to acknowledge the ANR for financial support. We would also like to thank G. Michel, B. Gallet, C. Touzé, and S. Varghese for fruitful discussions.
@@ Appendix A TABLES
$\eta_{\text{rms}(mm)}$00footnotetext: $\eta_{\text{rms}}$ is the r.m.s transversal displacement; $V_{\text{rms,inj}}$ is the r.m.s injected velocity; K is the kurtosis; S is the skewness; $1.159$ $1.158$ $2.230$ $3.596$
$V_{\text{rms,inj}(\frac{m}{s})}$ $0.179$ $0.180$ $0.312$ $0.464$
K00footnotetext: $N_{tot}$ is the total number of all recorded events; $N_{re}$ is the number of detected rare events; frac is the ratio of $\frac{N_{re}}{N_{tot}}$ $3.057$ $3.010$ $3.105$ $3.057$
@@
$N_{tot}$ $5.484\times 10^{5}$ $3.454\times 10^{6}$ $5.046\times 10^{5}$ $4.139\times 10^{5}$
$N_{re}$ $162$ $989$ $308$ $160$
frac $0.00029$ $0.00028$ $0.00061$ $0.00038$